# Measuring eccentricity and gas-induced perturbation from gravitational waves of LISA massive black hole binaries

### v1

@@ 7 Fitting a wrong template
To reduce statistical variance during PTMCMC, we take the average of Bayes factors from two sets of independent runs, while we report the recovered parameter from the first set of runs.
7.1 Injecting only either eccentricity or gas perturbation
@@
In Table 3, we show the recovered values of $e_{0}$ and Bayes factors when analysing an injected environmental circular signal for positive and negative values of $C_{\rm g}$ with an eccentric template. The stronger the negative $C_{\rm g}$, the higher the recovered eccentricity from its peaked Gaussian posterior. There are also larger biases on the intrinsic-merger parameters (see Appendix E). Since eccentricity can not explain an effect that slows down the inspiral caused by positive $C_{\rm g}$ (an outward torque), it is restricting $e_{0}$ to small values $\lesssim 10^{-3.5}$ with again large biases and one-sided Gaussian posteriors for eccentricity. For weak environmental effects, $C_{\rm g}\in\{-10^{2},-10^{1},10^{1}\}$, the Bayes factors are inconclusive. However, for strong environmental effects we can confidently accept the true template. Therefore, an EM counterpart will be crucial to be sure about the presence of a weak environmental effect.
$\boldsymbol{C_{g,\rm inj}}$ $\boldsymbol{\log_{10}e_{0}}$ $\boldsymbol{\ln\mathcal{B}}$ Strength
[…]
$-10^{1}$ $-2.6$ $-2.7$ Inconclusive
[…]
@@
Table 3: The injected environmental amplitude $C_{0,\rm inj}$, recovered $e_{0}$, Bayes factor $\ln\mathcal{B}$ and the strength ov evidence for the true model when fitting an eccentric vacuum model to a circular environmental signal. Here $\ln\mathcal{B}\gtrsim 5$ (green) represents definitive evidence in favor of the true model, $1\lesssim\ln\mathcal{B}\lesssim 3$ (lime) means the true model is weakly preferred, and $-3\lesssim\ln\mathcal{B}\lesssim-1$ (pink) implies the false model is weakly preferred (Taylor, 2021). We term $-3\lesssim\ln\mathcal{B}\lesssim 3$ as inconclusive and $\ln\mathcal{B}\gtrsim 5$ as decisive. We recover negligible eccentricities for positive migration ($C_{\rm g}>0$) since any significant $e_{0}$ cannot explain slower inspiral.
@@
In Table 4, we show the recovered value of $C_{\rm g}$ and Bayes factors, when fitting a gas perturbed circular template to an injected eccentric signal for various small eccentricities. We get narrow-peaked Gaussian posteriors for $C_{\rm g}$. Even for $e_{0,\rm inj}=10^{-2.5}$, we find Bayes factors in favor of the true model, which becomes stronger for higher eccentricity.
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ $\boldsymbol{C_{\rm g}}$ $\boldsymbol{\ln\mathcal{B}}$ Strength
[…]
@@
Table 4: Results obtained when analysing an injected eccentric vacuum signal with a circular environmental template. The columns show the injected eccentricity, $e_{0,\rm inj}$, recovered $C_{\rm g}$, Bayes factor, $\ln\mathcal{B}$, and the strength of evidence in favour of the true model.
@@ 7.2 Injecting both eccentricity and gas perturbation
In Table 5, we show the recovered value of $e_{0}$ when fitting a vacuum eccentric signal to an injected eccentric environmental signal. The posteriors for $e_{0}$ are narrow-peaked Gaussians. We find that for a strong environmental perturbation ($|C_{\rm g}|=10^{3}$) PTMCMC heavily favors the true model, while the results are inconclusive for ($|C_{\rm g}|=10^{2}$). This again emphasizes the need to have a complementary EM signal to have definite proof of an environment.
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}}$ $\boldsymbol{\log_{10}e_{0}}$ $\boldsymbol{\ln\mathcal{B}}$ Strength
[…]
$-2$ $-10^{2}$ $-1.9$ $-1.1$ Inconclusive
$-2$ $10^{2}$ $-2.2$ $-1.0$ Inconclusive
[…]
Table 5: The injected $e_{0,\rm inj}$ and $C_{0,\rm inj}$, recovered $e_{0}$, Bayes factor, and the strength of evidence in favour of the true model, when fitting an injected eccentric environmental signal with a vacuum eccentric template.
@@
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}}$ $\boldsymbol{C_{\rm g}}$ $\boldsymbol{\ln\mathcal{B}}$ Strength
[…]
@@ 8 Discussion
Most previous studies have assumed that near-merger MBHBs will be in a circular binary since the expectation is that the orbital eccentricity would be radiated away due to GW emission (Peters & Mathews, 1963; Peters, 1964) before the source reaches the LISA band ($e_{\rm LISA}\approx 0$). However, this has been challenged by recent suites of high-resolution hydrodynamical simulations that embed an eccentric MBHB in a thin accretion disc (Zrake et al., 2021; D’Orazio & Duffell, 2021; Siwek et al., 2023; Tiede & D’Orazio, 2024). These studies extrapolate to the LISA band using Peters (1964) orbital evolution equations to find $e_{\rm LISA}\sim 10^{-3}$ for the case of prograde discs, while $e_{\rm LISA}\sim 10^{-2}$ for retrograde CBDs given the stronger boosting of eccentricity. Similarly, the interaction with a stellar potential can lead to $e_{\rm LISA}\sim 10^{-3.5}$ (Gualandris et al., 2022). In recent work Garg et al. (2024), we found that a gas-induced eccentricity as low as $10^{-2.75}$ could be observable for MBHBs in a vacuum. As a consequence, neglecting eccentricity could lead to potential biases in the inference of the intrinsic-merger parameters, $\{M_{z},q,\chi_{1},\chi_{2},t_{c}\}$. These studies also emphasize that a gaseous environment can significantly alter the inspiral GW waveform of MBHBs not only via observable gas-induced dephasing (Garg et al., 2022; Dittmann et al., 2023), but also by exciting small but measurable eccentricities.
@@
Usually gas perturbations for MBHBs are measured in terms of a cumulative phase shift in the GW phase using Newtonian waveforms (Garg et al., 2022; Dittmann et al., 2023). While this is a good first step to estimate if gas could leave an observable imprint, it makes it impossible to relate this phase shift unequivocally to the presence of a CBD disc. This is because either higher PN corrections, a small eccentricity, or other environmental influences such as a dark matter spike or third body interaction can mimic this effect (Zwick et al., 2023). To be confident that this phase shift is most likely from gas, we need to measure both its effective amplitude ($C_{\rm g}$) and its power-law slope ($n_{\rm g}$). Since gas, via either migration or accretion, induces a phase correction at the $-4$PN order, we not only need high SNR but also numerous cycles in the inspiral phase to detect it. This is why EMRIs are traditionally preferred for this kind of study (e.g., by Speri et al. 2023), as a $q\gtrsim 10^{4}$ circular system spends $\sim 10^{5}$ GW cycles in band during a four-year LISA observation window, which allows for constraints on both $C_{\rm g}$ and $n_{\rm g}$ even for a moderate Eddington ratio of ${\rm f}_{\rm Edd}=0.1$. What MBHBs lack in the number of cycles, they make up for in their high SNRs. For this reason, we find that we would be able to confidently measure (relative error $<50$ per cent) $C_{\rm g}$ for a MBHB with $M_{z}=10^{6}~{}{\rm M}_{\sun}$ and $q=8$, if ${\rm f}_{\rm Edd}\sim 0.1$ for a circular system or ${\rm f}_{\rm Edd}\sim 1$ for an eccentric system, assuming a strong gravitational torque ($\xi\sim 10$) from an extremely thin accretion disc.
@@
Estimating the strength of a gas torque ($\xi$) on the MBHB near-merger is challenging, since most simulations study the system in the regime where GWs are not dominant. Current simulations for non-inspiraling MBHBs predict $\xi\lesssim 1$ for a moderately thin disc. There have been studies of the effects of gas on the GW phase of an EMRI in the LISA band (Derdzinski et al., 2019, 2021; Nouri & Janiuk, 2023), albeit using Newtonian-order approximations, but nothing yet for measuring gas torques on coalescing MBHBs aside from Dittmann et al. (2023). [^14] There are interesting studies on the relativistic accretion flows onto merging MBHBs which show complex accretion flows, however these are typically not evolved for long enough to measure accurate torques (see e.g. Avara et al. 2023; Ennoggi et al. 2023; Gutiérrez et al. 2022). Many of the simulations with MBHBs that focus on the long-term evolution of the binary neglect magnetic fields and radiative transfer, further increasing the uncertainty in the estimated torques on realistic systems. Therefore, it is not clear whether $\xi$ becomes stronger, weaker, or remains the same as a binary approaches merger. For this reason, in this study we have remained agnostic about the torque strength, and considered a wide-range of $\xi$ which allows for a $|C_{\rm g}|$ as high as $10^{3}$, in the case that future studies find stronger gas torques in the LISA regime.
@@
The effective gas-amplitude $C_{\rm g}$ in Eq. (12) depends on various binary-disc parameters including the accretion rate onto the binary. Our parameterization implies that a super-Eddington accretion rate may be responsible for a high value of $C_{\rm g}$, as well as a nonlinear dependence of the torque strength $\xi$ in the regime of extremely thin accretion disks which remains poorly explored. In other words, a degeneracy exists between the disc parameters (in addition to $f_{\rm Edd}$ and radiative efficiency) and the resulting torque ($\xi$). This can be broken by future simulations of these systems that explore more representative parameters for luminous AGN systems. Furthermore, inspiraling MBHBs are naturally expected to produce bright EM counterparts, which can provide valuable constraints on $f_{\rm Edd}$ and binary eccentricity. The presence of gas allows for a characteristic X-ray emission during the inspiral (see, e.g. Haiman, 2017; Dal Canton et al., 2019; Mangiagli et al., 2022; Cocchiararo et al., 2024). Detection of such counterparts will not only confirm the presence of an accretion disk, but also provide a narrower prior on the disc parameters from an independent channel. This would help to break the degeneracies between disc parameters such as ${\rm f}_{\rm Edd}$, $\xi$, and the radiative efficiency. The combination of this knowledge with measurements of environmental parameters from the GW signal will provide the strongest constraints on accretion disc structure and binary-disc interaction.
The concurrent measurement of gas-amplitude and initial eccentricity leads to worse constraints on both parameters. While it reduces the minimum measured $C_{\rm g}$ from $\sim 10^{2}$ for a circular orbit to $\sim 10^{3}$ for an eccentric system, it raises the eccentricity threshold from $e_{0}\sim 10^{-2.75}$ in vacuum to $\sim 10^{-2}$ in the presence of an accretion disc. Since $e_{\rm LISA}$ depends more on the disc parameters than the binary parameters for a near-equal mass MBHB (Zrake et al., 2021), we expect that higher ${\rm f}_{\rm Edd}$ or $\xi$ will lead to a higher eccentricity in the LISA band. Therefore, a concurrent measurement allows the possibility of measuring both inspiral effects to place, in-turn, stronger limits on the disc properties.
@@
We also find that a weak environmental perturbation ($|C_{\rm g}|\lesssim 10$) around a circular binary can be confused with an eccentric MBHB in a vacuum. This is not true for the vice versa case, as a vacuum eccentric signal can not be confused with a circular environmental system for small $e_{0}\gtrsim 10^{-2.5}$. This further emphasizes the importance of finding an EM counterpart to conclusively determine that an accretion disc is present. Otherwise, the parameters estimated from the inspiral part of the phase will include biases which can not be completely mitigated by information from the merger-ringdown part of the GW signal. A wrong analysis of the inspiral phase could thus lead to a false detection of eccentricity, and it could even possibly mimic a deviation from GR (Gair et al., 2013). This can have far-reaching consequences, such as raising doubts on the validity of GR, or leading to a biased distribution of MBHB parameters, which would be used to disentangle SMBH formation and growth channels.
@@
There are a number of caveats in the current work. The TaylorF2Ecc model does not include spin-eccentricity cross-terms in the phase, which should be negligible for the small eccentricities we study here. Moreover, we only consider the Newtonian GW amplitude without eccentric and gas-induced corrections. However, the inclusion of higher PN orders and eccentric-environmental cross-terms should only help to improve the measurements of the parameters. While the accuracy of TaylorF2Ecc reduces towards our cutoff at the innermost stable circular orbit, we expect the results will not change if an earlier cutoff is used (Garg et al., 2024). From the astrophysical perspective, the phase correction due to the gas perturbation given by Eq. (11) is a relatively simple model. It ignores any gas torque fluctuations during the binary orbit (Zwick et al., 2022) that, in the very early inspiral phase, produce a secular phase shift that mimics our model and can deviate from our fiducial CBD torque (=$\xi\dot{M}\Omega r^{2}$) even if torque fluctuations themselves are negligible after orbital average. Moreover, once the orbital averaged value of $\dot{a}_{\rm gas}$ does not well-approximate the semi-major axis decay rate towards the merger, these fluctuations can not be ignored. However, this time-domain effect is cumbersome to include consistently in the frequency-domain waveforms considered here, and we leave this to future work. Also, including higher PN terms in the gas phase would require even higher resolution hydrodynamical simulations so that we could go beyond the Newtonian equations of motion, whose rich features may help us find additional terms. Finally, we have always kept the radiative efficiency $\epsilon$ as $0.1$ even for our highly spinning MBHs. Even if the dependence of gas torque on $\epsilon$ remains the same near merger, it can vary for the same spin magnitude between prograde or retrograde. We have kept it fixed, and assume that any uncertainty is folded into $\xi$.

### v2

@@ 7 Fitting a wrong template
To reduce statistical variance during PTMCMC, we take the average of Bayes factors from two sets of independent runs, while we report the recovered parameter from the first set of runs. Moreover, we report errors on the Bayes factors by dividing the absolute difference between the Bayes factors from the independent runs by two. The estimated errors on $\ln\mathcal{B}$ suggest small variance between different Bayesian runs.
7.1 Injecting either only eccentricity or only gas perturbation
@@
In Table 3, we show the recovered values of $e_{0}$ and Bayes factors when analysing an injected environmental circular signal for positive and negative values of $C_{\rm g}$ with an eccentric template. The stronger the negative $C_{\rm g}$, the higher the recovered eccentricity from its peaked Gaussian posterior. There are also larger biases on the intrinsic-merger parameters (see Appendix E). Since eccentricity can not explain an effect that slows down the inspiral caused by positive $C_{\rm g}$ (an outward torque), it is restricting $e_{0}$ to small values $\lesssim 10^{-3.5}$ with again large biases and one-sided Gaussian posteriors for eccentricity. For weak environmental effects, $C_{\rm g}[10^{-15}]\in\{-10^{2},-10^{1},10^{1}\}$, the Bayes factors are inconclusive. However, for strong environmental effects we can confidently accept the true template. Therefore, an EM counterpart or a population-based inference (Garg et al., 2024a) will be crucial to be sure about the presence of a weak environmental effect. Furthermore, errors on Bayes factor do not change the conclusion about their decisiveness.
$\boldsymbol{C_{g,\rm inj}[10^{-15}]}$ $\boldsymbol{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log_{10}e_{0}}}$ $\boldsymbol{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ln\mathcal{B}}}$ Strength
$-10^{3}$ $-1.6$ $42.3^{+1.5}_{-1.5}$ Decisive
$-10^{2}$ $-2.0$ $3.5^{+0.7}_{-0.7}$ Inconclusive
$-10^{1}$ $-2.5$ $-2.3^{+0.2}_{-0.2}$ Inconclusive
$10^{1}$ $-4.4$ $1.1^{+0.1}_{-0.1}$ Inconclusive
$10^{2}$ $-3.8$ $80.8^{+0.2}_{-0.2}$ Decisive
$10^{3}$ $-4.0$ $310.8^{+0.3}_{-0.3}$ Decisive
@@
Table 3: The injected environmental amplitude $C_{g,\rm inj}$, recovered $e_{0}$, Bayes factor $\ln\mathcal{B}$ with errors, and the strength of evidence for the true model when fitting an eccentric vacuum model to a circular environmental signal. Here $\ln\mathcal{B}\gtrsim 5$ (green) represents definitive evidence in favor of the true model, $1\lesssim\ln\mathcal{B}\lesssim 3$ (lime) means the true model is weakly preferred, and $-3\lesssim\ln\mathcal{B}\lesssim-1$ (pink) implies the false model is weakly preferred (Taylor, 2021). We term $-3\lesssim\ln\mathcal{B}\lesssim 3$ as inconclusive and $\ln\mathcal{B}\gtrsim 5$ as decisive. We recover negligible eccentricities for positive migration ($C_{\rm g}>0$) since any significant $e_{0}$ cannot explain slower inspiral.
@@
In Table 4, we show the recovered value of $C_{\rm g}$ and Bayes factors, when fitting a gas perturbed circular template to an injected eccentric signal for various small eccentricities. We get narrow-peaked Gaussian posteriors for $C_{\rm g}$. Even for $e_{0,\rm inj}=10^{-2.5}$, we find Bayes factor in favor of the true model, which becomes stronger for higher eccentricity.
$\boldsymbol{\log_{10}e_{0,\rm inj}}$ $\boldsymbol{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{\rm g}}[10^{-15}]}$ $\boldsymbol{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ln\mathcal{B}}}$ Strength
$-2.5$ $-12.8$ $3.3^{+0.2}_{-0.2}$ Inconclusive
$-2.0$ $-121.1$ $5.3^{+0.2}_{-0.2}$ Decisive
$-1.5$ $-1441.9$ $58.0^{+0.1}_{-0.1}$ Decisive
@@
Table 4: Results obtained when analysing an injected eccentric vacuum signal with a circular environmental template. The columns show the injected eccentricity, $e_{0,\rm inj}$, recovered $C_{\rm g}$, Bayes factor $\ln\mathcal{B}$, and the strength of evidence in favour of the true model.
@@ 7.2 Injecting both eccentricity and gas perturbation
In Table 5, we show the recovered value of $e_{0}$ when fitting a vacuum eccentric signal to an injected eccentric environmental signal. The posteriors for $e_{0}$ are narrow-peaked Gaussians. We find that for a strong environmental perturbation ($|C_{\rm g}|=10^{-12}$) PTMCMC heavily favors the true model, while the results are inconclusive for ($|C_{\rm g}|=10^{-13}$). This again emphasizes the need to have a complementary EM signal to have definite proof of an environment.
$\boldsymbol{e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}[10^{-15}]}$ $\boldsymbol{\log_{10}e_{0}}$ $\boldsymbol{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ln\mathcal{B}}}$ Strength
$10^{-2}$ $-10^{3}$ $-1.6$ $40.6^{+0.3}_{-0.3}$ Decisive
$10^{-2}$ $-10^{2}$ $-1.9$ $1.2^{+1.2}_{-1.2}$ Inconclusive
$10^{-2}$ $10^{2}$ ${-2.4}$ $-1.2^{+0.6}_{-0.6}$ Inconclusive
$10^{-2}$ $10^{3}$ ${-2.3}$ $238.0^{+0.3}_{-0.3}$ Decisive
Table 5: The injected $e_{0,\rm inj}$ and $C_{g,\rm inj}$, recovered $e_{0}$, Bayes factor, and the strength of evidence in favour of the true model, when fitting an injected eccentric environmental signal with a vacuum eccentric template.
@@
$\boldsymbol{e_{0,\rm inj}}$ $\boldsymbol{C_{g,\rm inj}[10^{-15}]}$ $\boldsymbol{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{\rm g}}[10^{-15}]}$ $\boldsymbol{\ln\mathcal{B}}$ Strength
$10^{-2}$ $-10^{3}$ $-1122.6$ $2.6^{+0.7}_{-0.7}$ Inconclusive
$10^{-2}$ $-10^{2}$ $-220.5$ $1.7^{+1.1}_{-1.1}$ Inconclusive
$10^{-2}$ $10^{2}$ $-19.9$ $2.3^{+0.8}_{-0.8}$ Inconclusive
$10^{-2}$ $10^{3}$ $881.2$ $1.6^{+0.6}_{-0.6}$ Inconclusive
@@ 8 Discussion
The LISA Red Book (Colpi et al., 2024), which reflects the current science objectives of the community, does not consider eccentricity and gas effects in the analysis of future MBHB data. However, recent suites of high-resolution hydrodynamical simulations that embed an eccentric MBHB in a thin accretion disc (Zrake et al., 2021; D’Orazio & Duffell, 2021; Siwek et al., 2023; Tiede & D’Orazio, 2024) suggest measurable eccentricities $e_{\rm LISA}\gtrsim 10^{-2.75}$ (Garg et al., 2024b) in the LISA band despite partial circularization due to GW emission (Peters, 1964). These studies also emphasize that a gaseous environment can non-negligibly alter the inspiral GW waveform of MBHBs via observable gas-induced dephasing (Garg et al., 2022; Dittmann et al., 2023). Therefore, neglecting eccentricity and gas imprints in the GW waveform could induce bias as shown in Appendix E. This work can help to motivate the community to consider both gas and eccentricity as essential parameters for future data analysis.
@@
Usually gas perturbations for MBHBs are measured in terms of a cumulative phase shift in the GW phase using Newtonian waveforms (Garg et al., 2022; Dittmann et al., 2023). While this is a good first step to estimate if gas could leave an observable imprint, it makes it impossible to relate this phase shift unequivocally to the presence of a CBD disc. This is because either higher PN corrections, a small eccentricity, or other environmental influences such as a dark matter spike or third body interaction can mimic this effect (Zwick et al., 2023). To be confident that this phase shift is most likely from gas, we need to measure both its effective amplitude ($C_{\rm g}$) and its power-law slope ($n_{\rm g}$). Since gas, via either migration or accretion, induces a phase correction at the $-4$PN order, we not only need high SNR but also numerous cycles in the inspiral phase to detect it. This is why EMRIs are traditionally preferred for this kind of study (e.g., by Speri et al. 2023), as a $q\gtrsim 10^{4}$ circular system spends $\sim 10^{5}$ GW cycles in band during a four-year LISA observation window, which allows for constraints on both $C_{\rm g}$ and $n_{\rm g}$ even for a moderate Eddington ratio of ${\rm f}_{\rm Edd}=0.1$. However, what MBHBs lack in the number of cycles, they make up some of it by having high SNRs to allow us to measure disc properties within reasonable limits.
@@
Estimating the strength of a gas torque ($\xi$) on the MBHB near-merger is challenging, since most simulations study the system in the regime where GWs are not dominant. Current simulations for non-inspiraling MBHBs predict $\xi\lesssim 1$ for a moderately thin disc. There have been studies of the effects of gas on the GW phase of an EMRI in the LISA band (Derdzinski et al., 2019, 2021; Nouri & Janiuk, 2023), albeit using Newtonian-order approximations, but nothing yet for measuring gas torques on coalescing MBHBs aside from Dittmann et al. (2023). [^15] There are interesting studies on the relativistic accretion flows onto merging MBHBs which show complex accretion flows, however these are typically not evolved for long enough to measure accurate torques (see e.g. Avara et al. 2023; Ennoggi et al. 2023; Gutiérrez et al. 2022). Many of the simulations with MBHBs that focus on the long-term evolution of the binary neglect magnetic fields and radiative transfer, further increasing the uncertainty in the estimated torques on realistic systems. Therefore, it is not clear whether $\xi$ becomes stronger, weaker, or remains the same as a binary approaches merger. For this reason, in this study we have remained agnostic about the torque strength, and considered a wide-range of $\xi$ which allows for a $|C_{\rm g}|$ as high as $10^{-12}$, in the case that future studies find stronger gas torques in the LISA regime.
@@
Our parameterization in Eq. (18) implies that a super-Eddington accretion rate may be responsible for a high value of $C_{\rm g}$ as well as a nonlinear dependence of the torque strength $\xi$ in the regime of extremely thin accretion discs, which remains poorly explored. In other words, a degeneracy exists between the disc parameters (in addition to ${\rm f}_{\rm Edd}$ and radiative efficiency) and the resulting torque ($\xi$). This can be broken by future simulations of these systems that explore more representative parameters for luminous AGN systems. Furthermore, inspiraling MBHBs are naturally expected to produce bright EM counterparts, which can provide valuable constraints on ${\rm f}_{\rm Edd}$ and binary eccentricity. The presence of gas allows for a characteristic X-ray emission during the inspiral (see, e.g. Haiman, 2017; Dal Canton et al., 2019; Mangiagli et al., 2022; Cocchiararo et al., 2024). Detection of such counterparts will not only confirm the presence of an accretion disc, but also provide a narrower prior on the disc parameters from an independent channel. The observation of an EM counterpart will trigger the search for environmental deviations and justify the assumption of fixing $n_{\rm g}=4$, thus allowing us to break the degeneracies between disc parameters such as ${\rm f}_{\rm Edd}$, $\xi$, and the radiative efficiency. The combination of this knowledge with measurements of environmental parameters from the GW signal will provide the strongest constraints on accretion disc structure and binary-disc interaction.
@@
Section 7 emphasizes the importance of finding an EM counterpart or performing a population-based inference for moderate $C_{\rm g}$ to conclusively determine that an accretion disc is present. Otherwise, the parameters estimated from the inspiral part of the phase will include biases which can not be completely mitigated by information from the merger-ringdown part of the GW signal. A wrong analysis of the inspiral phase could thus lead to a false detection of eccentricity, and it could even possibly mimic a deviation from GR (Gair et al., 2013). This can have far-reaching consequences, such as raising doubts on the validity of GR, or leading to a biased distribution of MBHB parameters, which would be used to disentangle MBH formation and growth channels.
@@
There are a number of caveats in the current work. The TaylorF2Ecc model does not include spin-eccentricity cross-terms in the phase, which should be negligible for the small eccentricities we study here. Moreover, we only consider the Newtonian GW amplitude without eccentric and gas-induced corrections. However, the inclusion of higher PN orders and eccentric-environmental cross-terms should only help to improve the measurements of the parameters. While the accuracy of TaylorF2Ecc reduces towards our cutoff at the innermost stable circular orbit, we expect the results will not change if an earlier cutoff is used (Garg et al., 2024b). From the astrophysical perspective, the phase correction due to the gas perturbation given by Eq. (17) is a relatively simple model. It ignores any gas torque fluctuations during the binary orbit (Zwick et al., 2022) that, in the very early inspiral phase, produce a secular phase shift that mimics our model and can deviate from our fiducial CBD torque (=$\xi\dot{M}\Omega r^{2}$) even if torque fluctuations themselves are negligible after orbital average. Moreover, once the orbital averaged value of $\dot{a}_{\rm gas}$ does not well-approximate the semi-major axis decay rate towards the merger, these fluctuations can not be ignored. However, this time-domain effect is cumbersome to include consistently in the frequency-domain waveforms considered here, and we leave this to future work. Also, including higher PN terms in the gas phase would require even higher resolution hydrodynamical simulations so that we could go beyond the Newtonian equations of motion, whose rich features may help us find additional terms. Finally, we have always kept the radiative efficiency $\epsilon$ as $0.1$ even for our highly spinning MBHs. Even if the dependence of gas torque on $\epsilon$ remains the same near merger, it can vary for the same spin magnitude between prograde or retrograde. We have kept it fixed, and assume that any uncertainty is folded into $\xi$.